OpenStudy (blaque678):

Find the next 5 numbers. 1,5,12,22.....

OpenStudy (anonymous):

the sequence is, 1 1+(1+3)=5 5+(1+3+3)=12 12+(1+3+3+3)=22 22+(1+3*4)=35 35+(1+3*5)=51 51+(1+3*6)=70 70+(1+3*7)=92 92+(1+3*8)=117